In a certain Algebra 2 class of 29 students, 18 of them play basketball and 5 of them play baseball. There are 9 students who play neither sport. What is the probability that a student chosen randomly from the class plays basketball or baseball?


1
Expert's answer
2022-05-23T16:42:53-0400
"P(Basket\cup Base)=\dfrac{29-9}{29}=\dfrac{20}{29}"